OpenStudy (anonymous):

slove: 3x-4y=10 5x-3y=2

OpenStudy (anonymous):

x first by 3 and second by 4; 9x - 12y = 30 20x - 12y = 8 subtract -11x = 22 x = -2 plug x=-2 into first equn.: -6 - 4y = 10 4y = -16 y = -4 x=-2,y=-4